# Software Framework for Optically Accessible Quantum Memory Using Group-IV Color Centers in Diamond

### v2

@@ Abstract
In the rapidly evolving field of quantum technology, the precise and detailed description of quantum components is not just a necessity but the foundation for advancing research, development, and applications. Optically accessible quantum memories are key building blocks for devices such as quantum repeaters and two-factor authentication. The memory we describe here is based on a tin-vacancy color center coupled to a highly efficient cavity. It leverages state-dependent reflection from the cavity and implements high-fidelity fractional single qubit gates via a train of optical $\pi/8$ pulses. We also describe its operation under microwave control, further extending our analysis. Our primary contribution in this work is the integration of this device model into a standardized software framework for quantum memory architectures.
@@ I Introduction
Figure 1: a) A time-bin qubit $|\psi\rangle$ is stored in a Sawfish [1, 2] spin-photon interface that hosts a G4V. Storage is achieved via the electronic spin. Fast switches (S) are deployed throughout the circuit to enable efficient measurement and routing of the photonic state. The $X$-basis measurement is performed using a fast switch integrated with an imbalanced Mach-Zehnder interferometer. The storage procedure is completed once the photonic qubit is measured in the $X$-basis. For retrieval, the spin state is read out by entangling it with a photon from an extra single-photon source (SPS) after a time-bin preparation state (TBPS) [3, 4, 5]. A subsequent $Z$-basis measurement on the spin is performed. b) Using the SnV center as a G4V, spin-photon entanglement is mediated through a Sawfish cavity-to-fiber interface [1, 2]. The SnV is modeled as a four-level system characterized by its spontaneous emission rates $\gamma_{1A},\gamma_{2B}$, spectral contrast $\Delta\omega_{s}=\omega_{2B}-\omega_{1A}$ spin splitting $\omega_{s}$, and atomic transition frequencies $\omega_{1A},\omega_{2B}$, while interacting with a cavity mode $a_{c}$ at frequency $\omega_{c}$ for the spin-dependent reflection steps. c) Entanglement is generated employing a reflection scheme containing spin-dependent reflection steps and a $\pi/2$ rotation where the spin is initialized in the state $|1\rangle$. The $\pi/2$ rotation is applied to the spin before the final reflection event. This rotation can be implemented either using a sequence of four $\pi/8$ optical Raman pulses or utilizing microwave control. Figures b) and c) are adopted from [6].
@@
Optically accessible quantum memories become of increasing interest in the era of quantum communication [7, 8, 9], computing [10, 11, 12] and sensing [13, 14, 15]. Below, we highlight key applications: First generation quantum repeaters, a key component in scalable quantum networks, rely on a quantum memory to extend entanglement over long distances [16, 17, 18]. Beyond enabling long-distance entanglement, quantum memories also play a crucial role in quantum security applications [19] such as quantum authentication [20, 6]. While quantum memories are important for quantum security and communication they are also highly relevant for scalable quantum computing [21], and especially for the development of a quantum random-access-memory (QRAM) [22]. Finally, the field of quantum sensing benefits from the use of quantum memories since they can help in increasing the sensor’s sensitivity [23].
@@
There are various types of quantum memories: The atomic-ensemble based quantum memories can map a photonic qubit onto a collective-spin excitation shared between two atomic ensembles [24], or onto a spatially multiplexed cold atomic ensemble [25] achieving storage times in the range of tens of microseconds. Another type of memory is based on solid-state systems: Prominent examples are rare-earth ion-doped crystals [26] or color centers in diamond [27, 28], which are able to store quantum information using methods such as the atomic frequency comb (AFC) protocol [26, 29] or by using a spin-photon interface and long lived nuclear spins in the diamond lattice [27]. Superconducting quantum memories are also being actively researched [30, 31, 32, 33], which can be readily integrated into a superconducting quantum computing platform. Superconducting qubits are so far not optically accessible and require superconducting waveguides for networking [34]. Finally, hybrid quantum memories combine physically distinct systems to push the scheme’s performance beyond that of the individual components. Take for example storage and retrieval of photons from a quantum dot single-photon source interfaced with an atomic quantum memory [35], based on the Off-Resonant Cascaded Absorption (ORCA) protocol [36].
@@
In this work, we focus on a hybrid quantum memory scheme that uses single-photon sources [37], such as quantum dots [35], solid state emitter [38] or heralded single photons from spontaneous parametric down conversion [39], together with an atom-like memory system based on negatively charged group-IV color centers in diamond (G4V) [40]. The latter is integrated into a single-sided nanophotonic cavity, specifically the Sawfish cavity [1, 2]. Tin-vacancy centers (SnVs) are a particularly suitable choice for such quantum memories [28, 41, 42, 43] when integrated into nanophotonic structures. Due to their reduced sensitivity to charge noise [44], their optical coherence properties are largely preserved. By using state‑dependent reflection from the cavity and implementing high-fidelity fractional quantum gates through a pulse train of optical $\pi/8$ pulses, our specific memory design can achieve gate fidelities exceeding 99% depending on the bandwidth. We also consider microwave control for the memory scheme, which extends the viability to longer storage times, at the cost of reduced operational rates [45].
@@
The goal of this work is to present the technical background for describing the device within a standardized software framework for quantum memory architectures [46]. The software module we developed allows users to compute Kraus operators for the read-in and read-out processes under optical or microwave control, respectively, and to specify photon-generation fidelity, bandwidth, and the nanophotonic system’s temperature. When microwave control is selected, users may also adjust strain as well as DC and AC magnetic field strengths and orientations. To support this module, we provide a detailed description of the model for storing and retrieving a time-bin photonic qubit in a spin state, which serves as the basis of the software component.
@@
To model the spin–photon interface, we use closed-form integral expressions covering detailed modeling of spin-photon interaction for the read-in process, accounting for both an imperfect and broadband photon source and imperfect spin rotations, which are generated by optical control [6, 47] or by microwave control [45], depending on the color center’s temperature and strain. For the read-out process, we assume an ideal single-photon source but include imperfect spin rotations.
@@
Sec. II introduces our proposed hybrid quantum memory scheme consisting of a single-photon source and a group-IV color center in diamond. Sec. III outlines the mathematical principles concerning the described system. Sec. IV provides numerical examples for the Kraus-operators. Finally, Sec. V summarizes the results and provides an outlook.
@@ II Quantum Memory Scheme
The quantum memory scheme is illustrated in Figure 1. A single-photon source emits photons which are transmitted via a fiber to the quantum memory formed by a highly efficient single-sided Sawfish crystal cavity [1, 6] (see Figure 1b). These photons are entangled with the cavity’s spin and subsequently the photon is measured in the $X$-basis, completing the read-in process. For the read-out an extra single-photon source (SPS) is required. The respective photon from the SPS is entangled with the saved spin of the color center. Afterwards the spin is measured in the $Z$-basis. The reflection scheme for the read-in and read-out process contains an early reflection, a $\pi/2$ rotation, a late reflection and a measurement as illustrated in Figure 1a,c. The reflection steps realize a controlled phase gate which can be optimized by tuning the cavity loss rate $\kappa$, mode frequency $\omega_{c}$ and central frequency of the emitter $\omega_{0}$ depending on the bandwidth $\gamma$ of the photon. The spin rotation relies on coherent control of the electron spin of the SnV. That is done using optical [6] or microwave control [45]. Optical control is performed using a Raman scheme [6] applying a train of optical $\pi/8$ pulses to achieve a high fidelity $\pi/2$ rotation encountering photonic and phononic processes. The subsequent $X$-basis measurement and thereby the overall saved state after the read-in process represents an important set of the relevant error channels: depolarization during photon generation, broadband photons and control imperfections due to photonic and phononic processes inside the SnV [6]. Beyond [49] we take crosstalk during the spin-photon interaction into account.
@@
The read-out process uses an extra single-photon source whose photon gets entangled with the spin using the reflection scheme with a subsequent $Z$-measurement of the spin. In this work, we make the following key assumptions, which are fixed parameters in the memory software implementation.
@@
Single-photon source: We assume, that the SPSs photons have a fixed bandwidth and model the spin system as a four-level system. We assume a photon generation fidelity $F$, an incoming photonic qubit $|\psi\rangle=\alpha|e\rangle+\beta|l\rangle$ for some $(\alpha,\beta)$ with $|\alpha|^{2}+|\beta|^{2}=1$ and bandwidth $\gamma$. For the extra single-photon source we assume narrowband photons with $\alpha=\beta=1/\sqrt{2}$ with perfect fidelity, ensuring no limitations arising from the auxiliary photon source.
@@
Magnetic field: For the presented example calculation, we assume a static magnetic field strength $B_{\rm dc}=3.0$ T and magnetic field orientation $\theta_{\rm dc}=43.11\deg$ for performing optical control. In future iterations of our package we will include variable field strength and angle. For microwave control the static magnetic field strength $B_{\rm dc}$ as well as the magnitude of the microwave control field $B_{\rm ac}$ and their polar angles $\theta_{\rm dc}$ and $\theta_{\rm ac}$ are variable. We assume the azimuthal angles $\phi_{\rm dc}=0$ and $\phi_{\rm ac}=-\pi/2$ for achieving a rotation around the $y$-axis in the Bloch-sphere [45] (the orientations are visualized in Figure 1b.)
Strain: The SnV is in a low-strain environment for optical control and a strained one for microwave control with the compression $E_{x}$ and shear strain $\epsilon_{xy}$. These strain components can be freely chosen for microwave control.
@@
Temperature: The nanophotonic system including the SnV is assumed to be cryogenically cooled to temperatures ranging from $100$ mK to $4$ K.
Power: The average power consumption of the two lasers is calculated over the full-width at half maximum (FWHM) here,
$P_{i}^{L}=W_{i}/(\sigma\sqrt{e}),\quad W_{i}=1/2\epsilon_{0}E_{i}^{2}\lambda_{i}^{3}/(2n^{3})$ (1)
where $\sigma$ is the standard deviation of the laser pulse, $\epsilon_{0}$ is the dielectricity constant, $E_{i}$ for $i=1,2$ describes the laser field’s amplitude, $\lambda_{i}$ is the wavelength of the respective laser and $n$ is the refractive index of diamond. The microwave power is
@@
where $B_{\rm ac}$ is the magnetic field strength of the driving field, $A^{\rm MW}$ the effective area and $\mu_{0}$ the magnetic susceptibility.
Time: The total processing time is given by
@@
for $k=\rm read-in,read-out$ where $T_{\rm tb,k}$ is the time allocated for a time-bin qubit, $T_{g}$ is the control gate duration, $T_{m}$ is the measurement time, $T_{s}$ is the storage time, and $T_{c,k}$ is the transmission time for the respective process. We set $T_{\rm tb,k}=20T_{\rm lt,k}$, where $T_{\rm lt,k}$ is the lifetime of the single-photon source used for the read-in or read-out, respectively, and choose $T_{g}=40\sigma$ for optical control with $\sigma=\tau_{\pi/8}/2\sqrt{2\ln(2)}$, where $\tau_{\pi/8}=88.33$ ps is the full-width at half maximum of the optimized $\pi/8$ pulse for $B_{\rm dc}=3.0$ T. We take $T_{m}=100$ ps [50], corresponding to the dead time of the photon detectors [51]. The communication time is given by $T_{c,k}=L_{k}/c$, where $L_{k}$ is the fiber length for read-in or read-out, respectively and $c$ is the speed of light in the fiber.

### v3

# Software Framework for Optically Accessible Quantum Memories Using Group-IV Color Centers in Diamond

@@ Abstract
In the rapidly evolving field of quantum technology, the precise and detailed description of quantum components is not just a necessity but the foundation for advancing research, development, and applications. Optically accessible quantum memories are key building blocks for devices such as quantum repeaters and two-factor authentication. The memory we describe here is based on a Group-IV-vacancy color center (SiV-/SnV-) coupled to a highly efficient cavity. It leverages state-dependent reflection from the cavity and implements high-fidelity fractional single qubit gates via a train of optical $\pi/8$ pulses. We also describe its operation under microwave control, further extending our analysis. Our primary contribution in this work is the integration of this device model into a standardized software framework for quantum memory architectures.
@@ I Introduction
[…]
@@
Optically accessible quantum memories that can store and retrieve photonic qubits become of increasing interest in the era of quantum communication [7, 8, 9], computing [10, 11, 12] and sensing [13, 14, 15]. Below, we highlight key applications: First generation quantum repeaters, a key component in scalable quantum networks, rely on a quantum memory to extend entanglement over long distances [16, 17, 18]. Beyond enabling long-distance entanglement, quantum memories also play a crucial role in quantum security applications [19] such as quantum authentication [20, 6]. While quantum memories are important for quantum security and communication they are also highly relevant for scalable quantum computing [21], and especially for the development of a quantum random-access-memory (QRAM) [22]. Finally, the field of quantum sensing benefits from the use of quantum memories since they can help in increasing the sensor’s sensitivity [23].
@@
There are various types of quantum memories: The atomic-ensemble based quantum memories can map a photonic qubit onto a collective-spin excitation shared between two atomic ensembles [24], or onto a spatially multiplexed cold atomic ensemble [25] achieving storage times in the range of tens of microseconds. Another type of memory is based on solid-state systems: Prominent examples are rare-earth ion-doped crystals [26] or color centers in diamond [27, 28], which are able to store quantum information using methods such as the atomic frequency comb (AFC) protocol [26, 29] or by using a spin-photon interface and long lived nuclear spins in the diamond lattice [27]. Superconducting quantum memories are also being actively investigated [30, 31, 32, 33]. Although superconducting qubits have not yet been used to store individual photonic qubits, neither directly nor via transduction [34, 35], they have instead relied on superconducting waveguides for networking [36]. Recent progress has nevertheless been made in all-optical single-shot readout [37]. Finally, hybrid quantum memories combine physically distinct systems to push the scheme’s performance beyond that of the individual components. Take for example storage and retrieval of photons from a quantum dot single-photon source interfaced with an atomic quantum memory [38], based on the Off-Resonant Cascaded Absorption (ORCA) protocol [39].
@@
In this work, we focus on a hybrid quantum memory scheme that uses single-photon sources [40], such as quantum dots [38], solid state emitters [41] or heralded single photons from spontaneous parametric down conversion [42], together with an atom-like memory system based on negatively charged group-IV color centers in diamond (G4V) [43]. The latter is integrated into a single-sided nanophotonic cavity, specifically the Sawfish cavity due to its high coupling efficiency [1, 2]. Tin-vacancy centers (SnVs) are a particularly suitable choice for such quantum memories [28, 44, 45, 46] when integrated into nanophotonic structures. Due to their reduced sensitivity to charge noise [47], their optical coherence properties are largely preserved. By using state‑dependent reflection from the cavity and implementing high-fidelity fractional quantum gates through a pulse train of optical $\pi/8$ pulses, our specific memory design can achieve gate fidelities exceeding 99% depending on the bandwidth. We also consider microwave control for the memory scheme, which extends the viability to longer storage times, at the cost of reduced operational rates [48].
@@
The goal of this work is to present the technical background for describing the device within a standardized software framework for quantum memory architectures [49]. The software module we developed allows users to compute Kraus-operators for the read-in and read-out processes using the SiV- or SnV- as a quantum memory. The user can specify optical (limited for now to the SnV) or microwave control, the photon-generation fidelity, bandwidth for read-in and read-out, and the nanophotonic system’s temperature. When microwave control is selected, users may also adjust strain as well as DC and AC magnetic field strengths and orientations as well as the respective defect center, i.e. the SiV- or SnV-. If the user measures data for the spin gate or the cavity, these can be used to estimate the corresponding read-in and read-out Kraus-operators. To support this module, we provide a detailed description of the model for storing and retrieving a time-bin photonic qubit in a spin state, which serves as the basis of the software component.
@@
To model the spin–photon interface, we use closed-form integral expressions covering detailed modeling of spin-photon interaction for the read-in and read-out process, accounting for a broadband photon source and imperfect spin rotations, which are generated by optical control [6, 50] or by microwave control [48], depending on the color center’s temperature and strain.
@@
Sec. II introduces our proposed hybrid quantum memory scheme consisting of a single-photon source and a group-IV color center in diamond. Sec. III outlines the mathematical principles concerning the described system. Sec. IV demonstrates several use cases of the developed software. Finally, Sec. V summarizes the results and provides an outlook.
@@ II Quantum Memory Scheme
The quantum memory scheme is illustrated in Figure 1. A single-photon source (SPS 1) emits photons that are transmitted through a fiber to a quantum memory, implemented as a highly efficient single-sided Sawfish crystal cavity containing an SiV- or SnV- center [1, 6] (see Figure 1b). We chose the sawfish cavity as an example due to it large simulated efficiencies [1] exceeding $98\%$ efficiency of emitter to fiber coupling, but in principle any other half sided cavity spin system is admissible. These photons are entangled with the spin of the color center incorporated into the cavity and subsequently the photon is measured in the $X$-basis, completing the read-in process. For the read-out an extra single-photon source (SPS 2) is required. The respective photon from the SPS interacts with the spin-cavity system by combining the reflection scheme with a measurement in the $Z$-basis of the spin. The reflection scheme for the read-in and read-out process contains an early reflection, a $\pi/2$ rotation, a late reflection and a measurement as illustrated in Figure 1a,c. The reflection steps realize a controlled phase gate which can be optimized by tuning the cavity loss rate $\kappa$, mode frequency $\omega_{c}$ and central frequency of the emitter $\omega_{0}$ depending on the bandwidth $\gamma$ of the photon. The spin rotation relies on coherent control of the electron spin of the G4V. Control can be implemented using optical [6] or microwave control [48]. Optical control is performed using a Raman scheme [6] applying a train of optical $\pi/8$ pulses to achieve a high fidelity $\pi/2$ rotation in the presence of dissipative photonic and phononic processes. Notably, both microwave and optical spin control have been experimentally demonstrated for negatively charged silicon- and tin-vacancy (SiV-/SnV-) center in diamond [52, 53, 54, 55], highlighting the versatility of these platforms for quantum information applications. The subsequent $X$-basis measurement and thereby the overall saved state after the read-in process is affected by an important set of the relevant error channels: broadband photons and control imperfections due to photonic and phononic processes, which have an effect on the SiV-/SnV- [6, 48]. Beyond [56], we also account for crosstalk during the spin–photon interaction, originating from coupling of the cavity mode to off-resonant optical transitions (see Appendix, Sec. 1 for details).
@@
The read-out process uses an additional single-photon source (SPS 2) whose photon gets entangled with the spin using the reflection scheme in combination with a subsequent $Z$-measurement of the spin. In this work, we make the following key assumptions, which are fixed parameters in the memory software implementation.
@@
Single-photon source: We assume that the SPSs photons have a fixed bandwidth and model the spin system as a four-level system where the states are $|1\rangle,|2\rangle,|A\rangle,|B\rangle$. We assume an incoming photonic qubit $|\psi\rangle=\alpha|E\rangle+\beta|L\rangle$ for some $(\alpha,\beta)$ with $|\alpha|^{2}+|\beta|^{2}=1$ and bandwidth $\gamma_{\rm in}$. For the additional SPS we assume finite bandwidth photons with $\alpha=\beta=1/\sqrt{2}$, fidelity $F_{\rm aux}$ and bandwidth $\gamma_{\rm out}$.
@@
Magnetic field: For the presented example calculations, we assume static magnetic field strengths $B_{\rm dc}=0.3,1.0,3.0$ T and respective magnetic field orientations $\theta_{\rm dc}=81.4\deg\,,64.62\deg\,,43.11\deg$ from [6] for optical control of the SnV’s spin qubit. In future iterations of our package we will include variable field strength and angle as well as an optical control feature for the SiV. For microwave control the static magnetic field strength $B_{\rm dc}$ as well as the magnitude of the microwave control field $B_{\rm ac}$ and their polar angles $\theta_{\rm dc}$ and $\theta_{\rm ac}$ are variable. We assume the azimuthal angles $\phi_{\rm dc}=0$ and $\phi_{\rm ac}=-\pi/2$ for achieving a rotation around the $y$-axis in the Bloch-sphere [48] for both the SiV- and SnV- (the orientations are visualized in Figure 1b.)
Strain: The SnV- is in a low-strain environment for optical control [6]. The SiV-/SnV- is placed in a strained environment for microwave control, with the strain contribution given by $E_{x}=\epsilon_{xx}-\epsilon_{yy}$, where $\epsilon_{xx}$ and $\epsilon_{yy}$ are the normal strain components and $\epsilon_{xy}$ is the shear strain [48] component. The strain components can be independently adjusted to implement microwave control. The strain Hamiltonian is parameterized by a set of constants [48], which in the software module are represented by $(\alpha_{g},\alpha_{u},\beta_{g},\beta_{u})$. They are related to the physical strain components according to
$(\alpha_{g},\alpha_{u},\beta_{g},\beta_{u})=(d_{g}E_{x},\,d_{e}E_{x},\,2d_{g}\varepsilon_{xy},\,2d_{e}\varepsilon_{xy})\,$, (1)
where $d_{g}$ and $d_{e}$ are the ground- and excited-state strain coupling coefficients.
@@
Temperature: The nanophotonic system including the SiV-/SnV- is assumed to be cryogenically cooled to temperatures ranging from $100$ mK to $4$ K.
Power: For optical transitions, we normalize the laser powers by the saturation power and express them as the dimensionless ratio [57]
$s_{i}=\frac{2\Omega_{i}^{2}}{\gamma^{2}_{iA}+\Delta^{2}_{i}}$ (2)
where $\Omega_{i}=\mu_{i}E_{i}/\hbar$ is the Rabi-frequency with $E_{i}$ the amplitude of laser $i$, $\mu_{i}$ is the transition dipole moment coupling the levels $|i\rangle$ and $|A\rangle$ [6], $\gamma_{iA}$ is the spontaneous emission rate using Fermi’s golden rule from state $|A\rangle$ to $|i\rangle$ [50] and $\Delta_{i}$ is the detuning between the laser pulse $i$ and the excited state $|A\rangle$ [6]. The microwave power is
@@
where $B_{\rm ac}$ is the magnetic field strength of the driving field, $A^{\rm MW}$ the effective area with $\lambda^{\rm MW}$ the wavelength of the microwave pulse, $n$ the refractive index of diamond and $\mu_{0}$ the magnetic susceptibility.
Cooperativity: The cooperativity $C$ is a dimensionless parameter that quantifies the strength of light–matter interaction in a cavity QED system. It is defined as [58]
$C=\frac{g^{2}}{2\kappa\gamma_{a}},$ (4)
where $g$ is the single-photon coupling strength between the emitter and the cavity mode, $\kappa$ is the cavity field decay rate, and $\gamma_{a}$ is the atomic decay rate calculated from Fermi’s golden rule. Both $2\kappa$ and $\gamma_{a}$ are given as full widths at half maximum (FWHM) of the respective resonances. Since achieved cooperativities for SiV- cavities are limited to $25$ [59], and SnV- cavities reach only about $1.7$ [60], the user can define an upper bound for the cooperativity, which we denote by $C_{\rm max}$, defined by the enhancement of the transition from $|A\rangle$ to $|1\rangle$. Notably, the cooperativity between $|1\rangle$ and $|A\rangle$ is approximately as large as the cooperativity between $|2\rangle$ and $|B\rangle$ for G4Vs [56].
Our software package optimizes under cooperativity constraints. In addition, it allows users to specify their reported cavity parameters, since fabricated cavities often deviate from the optimized design. It can thus provide local optima useful for fabrication targets and estimate quantum memory performance for arbitrary cavity parameters and photon central frequencies.
Spin gate: For both microwave and optical control, a spin $\pi/2$ rotation is simulated if the user chooses to generate an optimized spin rotation. However, users may also specify the spin gate manually, for example when they are incorporating the routine into their own numerical simulations or interfacing it with experimental control hardware. In our module, users are required to provide the propagated states assuming the initial states
$\rho_{1}=|1\rangle\langle 1|,\quad\rho_{2}=|2\rangle\langle 2|,\quad\rho_{+}=|+\rangle\langle+|,\quad\rho_{R}=|R\rangle\langle R|$ (5)
with $|+\rangle=\frac{1}{\sqrt{2}}(|1\rangle+|2\rangle)$ and $|R\rangle=\frac{1}{\sqrt{2}}(|1\rangle+{\rm i}|2\rangle)$.
Time: The total processing time $T_{1}$ is given by
@@
for $k=\rm read-in,read-out$ where $T_{\rm tb,k}$ is the time allocated for a time-bin qubit, $T_{g}$ is the control gate duration, $T_{m}$ is the measurement time, $T_{s}$ is the storage time, and $T_{c,k}$ is the transmission time for the respective process. We set $T_{\rm tb,k}=20T_{\rm lt,k}$, where $T_{\rm lt,k}$ is the lifetime of the single-photon source used for the read-in or read-out, respectively, and choose $T_{g}=40\sigma$ for optical control with $\sigma=\tau_{\pi/8}/2\sqrt{2\ln(2)}$, where $\tau_{\pi/8}$ is the full-width at half maximum of the optimized $\pi/8$ pulse for the respective magnetic field strength listed in [6]. We take $T_{m}=100$ ps [61], corresponding to the dead time of the photon detectors [62]. The communication time is given by $T_{c,k}=L_{k}/c$, where $L_{k}$ is the fiber length for read-in or read-out, respectively and $c$ is the speed of light in the fiber.